Three 20-Ω resistors are connected in series to a 6-V battery that has only negligible internal resistance. What is the voltage difference across each resistor in V? (Hint: First calculate the current going through the three resistors and then find the voltage difference.)




1
Expert's answer
2021-12-07T09:51:53-0500

"R=R_1+R_2+R_3=20+20+20=60\ (\Omega)"


"I=V/R=6/60=0.1\ (A)"


"V_1=V_2=V_3=I\cdot R_1=0.1\cdot20=2\ (V)" . Answer